# Search for very-high-energy photon clusters and transients outside the galactic plane in the Fermi-LAT data

### v1

@@ 3.2 Transients
Blazars exhibit significant variability across their electromagnetic spectra, including at VHE part. While Fermi-LAT data have been used to study VHE variability in bright sources (Paliya, 2024), we focus on extreme, short-term variability (flares) in weaker sources where traditional light-curve analysis is impractical due to low photon counts. Such flares could manifest as spatially and temporally clustered VHE photons, where even doublets (two photons within days) can achieve high statistical significance given the low background at VHE. We identified 114 doublets of events separated by not more than 3 days (259200 seconds) and 0.3 degrees. Most of them originate from bright sources, such as Mrk 421 (57 doublets), Mrk 501 (16), 1ES 1959+650 (11), BL Lac (5), 1ES 1218+304 (4), PKS 2155-304 (4), PG 1553+113 (3), 3FHL J0508.0+6737333A designation in the 3FHL catalog is given when there is no TeVCat association (2), 1ES 1011+496 (1), 1ES 1727+502 (1), 3FHL J1457.8-4642 (1), 3FHL J0308.4+0408 (1), 3FHL J0847.0-2337 (1), 3FHL J1203.3-3924 (1), 3FHL J1037.6+5711 (1), 3C 66A (1), PKS 1440-389 (1), PKS 0447-439 (1), S5 0716+714 (1), and one orphan flare without counterparts in 4FGL, 3FHL and TeVCat catalogs. Additionally, we detected 12 triplets (10 from Mrk 421, 2 from 1ES 1959+650) and 1 quadruplet (Mrk 421). We focused on 6 doublets from sources without TeVCat associations or prior VHE flare reports (excluding 3FHL J0508.0+6737, which exhibited a VERITAS-detected flare in 2009 (Ong, 2009)). These doublets show exceptional significance ($P_{glob}\in(10^{-8},~{}10^{-4})$), even without accounting for spatial coincidence with 3FHL sources (Table 4).
Given the low statistics, two photons constrain the flare’s expected event count to $6.50.35<N_{flare}<$ 6.5 (95% CL). This translates to VHE flux limits: $0.17F_{VHE}^{0}<F_{VHE}<3.25F_{VHE}^{0}$, where $F_{VHE}^{0}=2/\mathcal{E}$ and $\mathcal{E}$ is the Fermi-LAT exposure during the doublet’s time window. For each candidate flare, we analyzed all VHE photons within $0.3^{\circ}$ of the associated source and and 10-100 GeV photons [^4] We will refer to them as HE photons in the following. in this circle in time interval ($t_{0}-10^{7}$ s, $t_{0}+10^{7}$ s), where $t_{0}$ is the doublet’s midpoint.
@@
The VHE and HE photon lists for this source are provided in Appendix Tables 5 and 6. Only two VHE photons were detected within $0.3^{\circ}$ of the source over 16 years, implying a flare with a flux increase of $>10^{3}$ relative to the quiescent VHE level. Notably, only one of 36 HE photons coincided with the flare epoch (observed 5.5 Ms prior to the VHE event), suggesting an orphan VHE flare without a HE counterpart.
NGC 1218 is an S0 galaxy hosting the radio-bright Seyfert Type I nucleus 3C 78. While Seyfert galaxies sometimes exhibit GeV flaring activity (Gokus et al., 2021), VHE flares have been reported only in the Type II Seyfert NGC 1275 (Perseus Cluster core; (MAGIC Collaboration et al., 2018; Cao et al., 2024)). Intriguingly, recent IceCube observations associate Seyferts like NGC 1068 and NGC 7469 with astrophysical neutrinos (IceCube Collaboration et al., 2022; Sommani et al., 2025), raising questions about their potential as transient multi-messenger sources.
@@
PMN J0847-2337 is a candidate blazar exhibiting an unusual temporal pattern: a GeV-flare detected after a VHE one. Analysis of its photon distribution (Tables 10, 11) reveals that 8 of 90 HE photons arrived within a 3 Ms window starting 4 Ms after the VHE flare, with 5 photons concentrated in a 10-day sub-interval. Naturally, there is a strong observational bias against detection of such flares with IACTs, which usually exploit HE flares as a sort of external triggers.
@@
This relatively bright BL Lac object exhibits 10 VHE and 325 HE photons in total. The VHE flare coincides with a prolonged period of enhanced HE activity (Tables 13, 14), and a transient event was independently detected by HAWC approximately one week after the Fermi-LAT doublet (Weisgarber and HAWC Collaboration, 2018).
Despite its brightness, this source lacks a secure redshift measurement (Table 4). Spectroscopic observations in SDSS DR7 reported $z=0.83\pm 0.0004$ but with low confidence ($z_{conf}=0.07$) (Abazajian et al., 2009). Follow-up observations with larger telescopes revealed a featureless optical spectrum, yielding only a lower limit of $z>0.25$ (Paiano et al., 2020; Kasai et al., 2023). Photometric estimates from broadband SED modeling gave a less reliable but intriguing value of $z=1.14$ (Foschini et al., 2022).
The source’s AGN-dominated emission ($\sim 0.5$ mag brightening over two decades) further obscures host galaxy features, complicating redshift determination.
@@
This source is a blazar candidate, possibly of BL Lac type. It exhibited VHE flare without HE counterpart, and none of 40 HE photons arrived in 20 Ms interval of interest (see Table 9).
@@
Only two VHE photons were detected from this blazar candidate during the entire 16-year time span of the observations, both constituting a temporally clustered doublet. This VHE activity was preceded by enhanced HE emission: 4 out of 12 HE photons arrived within a 1.5 Ms window starting 3 Ms prior to the VHE event (Tables 7, 8).
This HE precursor to VHE flaring is a common feature of blazar variability. For instance, during the 2011 outburst of BL Lacertae, HE activity preceded VHE emission by $\sim$ month (Arlen et al., 2013).
@@ 3.2.6 NGC 5549?
The lack of a 3FHL/4FGL counterpart complicates the identification of a host source for this VHE doublet. While several galaxies reside within the same galaxy group as NGC 5549 – a low-luminosity AGN exhibiting faint nuclear activity – NGC 5549 itself emerges as the most plausible candidate due to its weak but discernible AGN signatures.
@@ 3.2.7 Energetics and spectral properties of the flares
Total absence of HE emission could pose a challenge to standard synchrotron self-Compton (SSC) model, suggesting that some alternative mechanisms of emission could be responsible for such flares.
@@ 4 Summary and conclusions
In this paper we searched for spatio-temporal clusters of the VHE events ($E>100$ GeV) in 16 years of the Fermi-LAT data at galactic latitudes $|b|>10^{\circ}$. Using DBSCAN algorithm, we found 107 clusters. All, but three of them, were associated with already known VHE sources, listed in 3FHL or TeVCat catalogs (Table LABEL:tab:clusters_full_list). From remaining three (Tables 1, 2), 4FGL J1544.3-0649, which had an epoch of activity in 2017-2018, could belong to a rare class of transient blazars and because of that evaded inclusion into VHE catalogs. Other sources, 4FGL J0039.1-2219 and 4FGL J0039.1-2219 are weak BL Lacs, that could potentially be detected with the IACTs.
We also searched for flaring VHE activity, looking for event multiplets within short temporal intervals of 3 days. We found 114 doublets, most of them (57) from Mrk 421 – the brightest extragalactic source with a very high level of variability. We focused on 6 doublets without TeVCat sources in their proximity. Five of them were associated with 3FHL sources (J0308.4+0408, J0847.0-2337, J1037.6+5711, J1037.6+5711, J1457.8-4642), and one has no association with any known HE/VHE sources (see Table 4).
Due to a limited effective area of the Fermi-LAT this approach is sensitive only to the brightest flares with a very high luminosity that in cases of 3FHL J1037.6+5711 and 3FHL J1457.8-4642 could reach $\mathcal{O}(10^{47}~{}\mathrm{erg~{}s^{-1}})$ (isotropic), which is, e.g., order of magnitude larger than the brightest BL Lac flare in July 2021.
One of the sources – 3FHL J0308.4+0408(NGC 1218) – is a Seyfert Type I galaxy. No objects of this type were known VHE sources before. A fortiori, no short energetic VHE flares was anticipated from such objects. Even more striking is plausible association of ’orphan’ doublet with LINER galaxy NGC 5549. This type of galaxies are considered to harbor low-luminosity AGNs in their centers and currently they are below sensitivity limit of HE/VHE telescopes. Still, there was a major extremely hard flare, during which the VHE flux ($E>100$ GeV) exceeded the average flux at HE ($E>10$ GeV) at least thousandfold, and the VHE luminosity of the AGN surpassed $10^{45}~{}\mathrm{erg~{}s^{-1}}$. It is worth noting, that non-blazar AGNs are a small minority in full population of TeV sources, with only 4 Fanaroff-Riley I galaxies are present in the TeVCat (M87, 3C 264, NGC 1275, and Cen A), so a detection of VHE flares from low-luminous non-aligned AGNs could have certain significance for a field of AGN jet studies.
@@
Note: When this paper was in a final stage of preparation, another work in which DBSCAN algorithm was used to search for clusters in GeV Fermi data, appeared (Cozzolongo et al., 2025). Their analysis focused on identifying extended sources in the Galactic plane, contrasting with our emphasis on extragalactic transients and high-latitude sources.
@@ Appendix A Table of clusters
Appendix B Table of transients
Table 4: Properties of sources associated with transient doublet events. Columns show: (1) 3FHL designation; (2) angular separation from doublet center (degrees); (3) common source name; (4) redshift from (Foschini et al., 2022) compilation (boldface indicates spectroscopic measurements, regular font photometric estimates); (5) source type (BLL = BL Lacertae, Sy 1 = Seyfert Type I, LIN = LINER galaxy); (6) 1–100 GeV flux from 4FGL in $10^{-9}~{}\mathrm{cm^{-2}~{}s^{-1}}$; (7) global p-value for transient detection; (8) doublet duration $\Delta t=t_{2}-t_{1}$ in $10^{5}~{}\mathrm{s}$; (9) Fermi-LAT exposure time in $10^{8}~{}\mathrm{cm^{2}~{}s}$; (10) observed VHE flux $F_{VHE}^{0}$ with 95% CL (0.17–3.3)×$\times F_{VHE}^{0}$ in $10^{-9}~{}\mathrm{cm^{-2}~{}s^{-1}}$; (11) derived luminosity in $10^{45}~{}\mathrm{erg~{}s^{-1}}$
@@ Appendix B Table of transients
Appendix C Lists of events for transients
@@
$t$, s $E$, GeV $l$, $b$, Distance,
580411826 156.0 353.234 61.364 0.04
580628550 592.9 353.25 61.437 0.03
Table 12: List of VHE photons for NGC5549(?), analogous to Table 5

### v2

@@ 3.2 Transients
Blazars exhibit significant variability across their electromagnetic spectra, also in the VHE range. While Fermi-LAT data have been used to study VHE variability in bright sources (Paliya, 2024), we focus on extreme, short-term variability (flares) in weaker sources where traditional light-curve analysis is impractical due to low photon counts. Such flares could manifest as spatially and temporally clustered VHE photons, where even doublets (two photons within a few days) can achieve high statistical significance given the low background at VHE. We identified 114 doublets of events separated by not more than 3 days (259200 seconds) and 0.3 degrees. Most of them originate from bright sources, such as Mrk 421 (57 doublets), Mrk 501 (16), 1ES 1959+650 (11), BL Lac (5), 1ES 1218+304 (4), PKS 2155-304 (4), PG 1553+113 (3), 3FHL J0508.0+6737333A designation in the 3FHL catalog is given when there is no TeVCat association (2), 1ES 1011+496 (1), 1ES 1727+502 (1), 3FHL J1457.8-4642 (1), 3FHL J0308.4+0408 (1), 3FHL J0847.0-2337 (1), 3FHL J1203.3-3924 (1), 3FHL J1037.6+5711 (1), 3C 66A (1), PKS 1440-389 (1), PKS 0447-439 (1), S5 0716+714 (1), and one orphan flare without counterparts in the 4FGL, 3FHL and TeVCat catalogs. Additionally, we detected 12 triplets (10 from Mrk 421, 2 from 1ES 1959+650) and 1 quadruplet (Mrk 421). We focused on the 6 doublets from sources without TeVCat associations or prior VHE flare reports (excluding 3FHL J0508.0+6737, which exhibited a VERITAS-detected flare in 2009 (Ong, 2009)). These doublets show exceptional significance ($P_{\mathrm{glob}}\in(10^{-8},~10^{-4})$), even without accounting for the spatial coincidence with 3FHL sources (Table 3).
Given the low statistics, two photons constrain the flare’s expected event count to $0.35<N_{flare}<6.5$ (95% CL). This translates to the VHE flux limits: $0.17F_{VHE}^{0}<F_{VHE}<3.25F_{VHE}^{0}$, where $F_{VHE}^{0}=2/\mathcal{E}$ and $\mathcal{E}$ is the Fermi-LAT exposure during the doublet’s time window. For each candidate flare, we analyzed all VHE photons within $0.3^{\circ}$ of the associated source and and 10-100 GeV photons [^4] We will refer to them as HE photons in the following. in this circle in the time interval ($t_{0}-10^{7}$ s, $t_{0}+10^{7}$ s), where $t_{0}$ is the doublet’s midpoint.
@@
The VHE and HE photon lists for this source are provided in Appendix Tables 4 and 5. Only two VHE photons were detected within $0.3^{\circ}$ of the source over 16 years, implying a flare with a flux increase of $10^{3}$ relative to the quiescent VHE level. Notably, only one of 36 HE photons coincided with the flare epoch (observed 5.5 Ms prior to the VHE event), suggesting an orphan VHE flare without a HE counterpart.
NGC 1218 is a S0 galaxy hosting the radio-bright Seyfert Type I nucleus 3C 78. While Seyfert galaxies sometimes exhibit GeV flaring activity (Gokus et al., 2021), VHE flares have been reported only in the Type II Seyfert NGC 1275 (Perseus Cluster core; (MAGIC Collaboration et al., 2018; Cao et al., 2024)). Intriguingly, recent IceCube observations associate Seyferts like NGC 1068 and NGC 7469 with astrophysical neutrinos (IceCube Collaboration et al., 2022; Sommani et al., 2025), raising questions about their potential as transient multi-messenger sources.
@@
PMN J0847-2337 is a candidate blazar exhibiting an unusual temporal pattern: a GeV-flare detected after a VHE one. Analysis of its photon distribution (Tables 6, 7) reveals that 8 of 90 HE photons arrived within a 3 Ms window starting 4 Ms after the VHE flare, with 5 photons concentrated in a 10-day sub-interval. Naturally, there is a strong observational bias against detection of such flares with IACTs, which usually exploit HE flares as a sort of external triggers.
@@
This relatively bright BL Lac object exhibits 10 VHE and 325 HE photons in total. The VHE flare coincides with a prolonged period of enhanced HE activity (Tables 8, 9), and a transient event was independently detected by HAWC approximately one week after the Fermi-LAT doublet (Weisgarber and HAWC Collaboration, 2018).
Despite its brightness, this source lacks a secure redshift measurement (Table 3). Spectroscopic observations in SDSS DR7 reported $z=0.83\pm 0.0004$ but with low confidence ($z_{\mathrm{conf}}=0.07$) (Abazajian et al., 2009). Follow-up observations with larger telescopes revealed a featureless optical spectrum, yielding only a lower limit of $z>0.25$ (Paiano et al., 2020; Kasai et al., 2023). Photometric estimates from broadband SED modeling gave a less reliable but intriguing value of $z=1.14$ (Foschini et al., 2022).
The source’s AGN-dominated emission ($\sim 0.5$ mag brightening over two decades) further obscures host galaxy features, complicating the redshift determination.
@@
This source is a blazar candidate, possibly of the BL Lac type. It exhibited a VHE flare without a HE counterpart, and none of the 40 HE photons arrived in the 20 Ms interval of interest (see Table 10).
@@
Only two VHE photons were detected from this blazar candidate during the entire 16-year time span of the observations, both constituting a temporally clustered doublet. This VHE activity was preceded by enhanced HE emission: 4 out of the 12 HE photons arrived within a 1.5 Ms window starting 3 Ms prior to the VHE event (Tables 11, 12).
This HE precursor to VHE flaring is a common feature of blazar variability. For instance, during the 2011 outburst of the BL Lacertae, HE activity preceded VHE emission by $\sim$ month (Arlen et al., 2013).
@@ 3.2.6 NGC 5549?
The lack of a 3FHL/4FGL counterpart complicates the identification of the host source for this VHE doublet. While several galaxies reside within the same galaxy group as NGC 5549 – a low-luminosity AGN exhibiting faint nuclear activity – NGC 5549 itself emerges as the most plausible candidate due to its weak but discernible AGN signatures.
@@ 3.2.7 Energetics and spectral properties of the flares
The total absence of HE emission could pose a challenge to the standard synchrotron self-Compton (SSC) model, suggesting that some alternative mechanisms of emission could be responsible for such flares.
Detected flares bear striking resemblance to several instances of fast variability of another misaligned sources, i.e. IC 310 and M 87 (Rieger and Levinson, 2018; Rieger, 2019): in both cases we need some new additional VHE component to account for the observed spectrum. Still, if we took quiescent HE luminosity as a proxy for a jet power $L_{j}$ (Nemmen et al., 2012), VHE -jet luminosity ratios $L_{\mathrm{VHE}}/L_{j}$ have considerably higher values for our set, which possibly disfavor models with magnetospheric origin of the emission and various jet-object interaction scenarios.
@@ 4 Summary and conclusions
In this paper we searched for spatio-temporal clusters of the VHE events ($E>100$ GeV) in 16 years of the Fermi-LAT data at galactic latitudes $|b|>10^{\circ}$. Using the DBSCAN algorithm, we found 107 clusters. All, but three of them, were associated with already known VHE sources, listed in the 3FHL or TeVCat catalogs (Table LABEL:tab:clusters_full_list). From remaining three (Table 2), 4FGL J1544.3-0649, which had an epoch of activity in 2017-2018, could belong to a rare class of transient blazars and because of that evaded inclusion into VHE catalogs. The other sources, 4FGL J0039.1-2219 and 4FGL J0039.1-2219 are weak BL Lacs, that could be a tip of an iceberg of much wider population of weaker sources, that would be detected with the next-generation IACTs.
We also searched for flaring VHE activity, looking for event multiplets within short temporal intervals of 3 days. We found 114 doublets, most of them (57) from Mrk 421 – the brightest extragalactic source with a very high level of variability. We focused on the 6 doublets without TeVCat sources in their proximity. Five of them were associated with 3FHL sources (J0308.4+0408, J0847.0-2337, J1037.6+5711, J1203.3-3924, J1457.8-4642), and one has no association with any known HE/VHE sources (see Table 3).
Due to the limited effective area of the Fermi-LAT this approach is sensitive only to the brightest flares with a very high luminosity that in cases of 3FHL J1037.6+5711 and 3FHL J1457.8-4642 could reach $\mathcal{O}(10^{47}~\mathrm{erg~s^{-1}})$ (isotropic), which is, e.g., an order of magnitude larger than the brightest BL Lac flare in July 2021.
One of the sources – 3FHL J0308.4+0408(NGC 1218) – is a Seyfert Type I galaxy. No objects of this type not belonging to narrow-line Sy 1 subclass were known as VHE sources before. A fortiori, no short energetic VHE flares were anticipated from such objects. Even more striking is the plausible association of the ’orphan’ doublet with the LINER galaxy NGC 5549. This type of galaxies are considered to harbor low-luminosity AGNs in their centers and currently they are below sensitivity limit of HE/VHE telescopes. Still, there was a major extremely hard flare, during which the VHE flux ($E>100$ GeV) exceeded the average flux at HE ($E>10$ GeV) at least a thousandfold, and the VHE luminosity of the AGN surpassed $10^{45}~\mathrm{erg~s^{-1}}$. It is worth noting, that non-blazar AGNs are a small minority in the full population of TeV sources, with only 4 Fanaroff-Riley I galaxies present in the TeVCat (M87, 3C 264, NGC 1275, and Cen A), so a detection of VHE flares from low-luminous non-aligned AGNs could have considerable significance for the field of AGN jet studies.
@@
Note: When this paper was in a final stage of preparation, another work in which a DBSCAN algorithm was used to search for clusters in GeV Fermi data, appeared (Cozzolongo et al., 2025). Their analysis focused on identifying extended sources in the Galactic plane, contrasting with our emphasis on extragalactic transients and high-latitude sources.
@@ Appendix A Table of clusters
Appendix B Table of VHE cluster candidates
# $l,^{\circ}$ $b,^{\circ}$ $N_{\mathrm{ph}}$ $P_{\mathrm{glob}}$ 4FGL name $F_{1-100}$ $\gamma$ $z$
93 93.36 -84.35 5 $7.9\times 10^{-7}$ J0039.1-2219 2.1 1.8 0.064
101 0.30 36.15 7 $2.5\times 10^{-12}$ J1544.3-0649 12.1 1.9 0.17
102 164.34 -58.59 5 $3\times 10^{-6}$ J0212.2-0219 3.1 2.0 0.25
Table 2: Properties of unassociated VHE cluster candidates and their associations in 4FGL. Columns: (1) Cluster ID from Table LABEL:tab:clusters_full_list; (2,3) Galactic longitude and latitude of cluster centroid; (4) Photon count; (5) Global probability of chance occurrence; (6) Associated 4FGL source; (7) Integrated GeV flux from 4FGL in $10^{-10}~\mathrm{cm^{2}~s^{-1}}$; (8) Spectral index; (9) Redshift.
Appendix C Table of transients
@@
3FHL name $\rho$, Assoc. name $z$ type $F_{1-100}$ $P_{\text{glob}}$ $\Delta t$ $\mathcal{E}$ $F_{VHE}^{0}$ $L$
@@
Appendix D Lists of events for transients
@@
$t$, s $E$, GeV $l$, $b$, Distance,
260094483 101.2 293.114 22.563 0.21
373271798 172.2 292.963 22.533 0.07
373409583 203.3 292.754 22.634 0.19
694676662 101.9 292.877 22.501 0.02
$t$, s $E$, GeV $l$, $b$, Distance,
391732356 158.7 324.526 10.815 0.15
391881102 148.2 324.357 11.010 0.16
$t$, s $E$, GeV $l$, $b$, Distance,
388126965 17.7 324.369 10.825 0.02
388309964 55.8 324.365 10.840 0.01
389170862 30.9 324.383 10.801 0.05
389618335 10.8 324.459 10.743 0.13
$t$, s $E$, GeV $l$, $b$, Distance,
580411826 156.0 353.234 61.364 0.04
580628550 592.9 353.25 61.437 0.03
Table 13: List of VHE photons for NGC5549(?), analogous to Table 4